OpenStudy (anonymous):

There are 9 blue marbles,4 black marbles,5 white marbles, and 6 red marbles in a box. Six marbles are drawn at random.If the probability of drawing a blue marble is now 1/3 ,how many of the 6 marbles drawn were blue?

OpenStudy (deoxna):

The total marbles are 24 (9+4+5+6=24) So the probability of taking out a blue one is 9/24 You take 6 marbles from the total so the fraction would be 9-x/18 (24-6=18), where x is the number of blue marbles taken out. So then 9-x/18=1/3, do the math and x=3 (6/18=1/3). The number of blue marbles drawn out was 3.

OpenStudy (anonymous):

after drawing 6 marbles there are 18 left so for probability to be 1/3 tyere must be 6 marbles left - so 3 were drawn

OpenStudy (anonymous):

3 blue marbles were drawn out

OpenStudy (anonymous):

Thanks Guys